Week of 060904

Open Actions from last week:
  • CMS see bursts of errors (still continuing monday morning). - Michael to send timestamps, Miguel to investigate. (low priority)
  • VOMS102 - why no automated procedure for restart ? (Harry)
  • New version lcg-gridftp-mon with cronjob (Maarten) - PENDING till next week
  • Some glite rpms will be upgraded today on grid services (mostly R-GMA, edg-mkgridmap) (Jan) - SLC4 DONE, SLC3 on monday - DONE
  • Monitoring of essential daemons on RB needed (GMOD, James)
  • FTS - disable atuators when going into maintenance so we don't have to turn off monitoring (Gavin) -DONE
  • 4AM Gridftp alarms - use new lemon features to smooth out these transient regular alarms (Jan) -DONE
  • Alice VOBOX to be replaced today (lxb2065) - Ulrich -DONE
  • CMS have asked for 2 'rb102's for production usage. Will need 1 new machine (prob from ALICE) and a WN for the test service currently on rb102. (Harry) - IN PROGRESS

Chair: James

SMOD: Thorsten

GMOD: David

Monday:

Log: FTA_WRONG on fts101 @ 4AM. rb108 - FS full, gdrb01-06, alarms on sure.

New Actions:

  • Track progress of remedy ticket for FS Full on rb108 to see if new proc(send alarms to remedy) works (James) - DONE
  • gridftp alarms still there on two nodes after monitoring reconfig - check. (Jan) IN PROGRESS
  • Followup GSSDATLAS reported problems with CASTOR transfers (Olof) DONE
  • Short queues created on CEs. Reconfig today at 12. Transparent for LCG CEs, short intervention for gLite CEs (ulrich) - DONE
  • Reminders to be send/announced for C2SC4 cluster shutdown new week for reconfig. no dteam/ops transfers to CERN all next week (Jan) - DONE

Discussion:

  • CMS are still preparing for tape transfer tests. Will start with PIC/Lyon. the tt-bar sample was successfully imported into CERN.
  • GSSDATLAS have a reported problem with transfers into CASTOR. To be followed up. Also AFS software area issues (too many accesses)

Tuesday:

Log:RAS

New Actions:

Discussion:

  • Identified replacement test rb (lxb7026) and Bernd has a spare node suitable for the new production RB. It's in the critical area, and will need moved.
  • GSSDATLAS user created 200K small files in CASTOR. User contacted. Zhongliang to also follow up.

Wednesday

Log: RAS

Actions:

Discussion:

  • upgrade to gLite 3.0.2 to be done by all service managers
  • CAs are currently written into shared area in AFS by GD user (flavia). This should be done in a more controlled manner. To be followed up in LCG SCM

Thursday

Log: CERN VACATION

Actions:

Discussion:

Friday

Log:

Actions:

  • Shutdown C2SC4 Cluster on monday for 1 week (Jan)

Discussion:

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2007-02-02 - FlaviaDonno
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback